bhalu corporation, an all equity, no growth firm recently hired an Alliant MBA as a financial analyst in their Finance Department. she proposed that Bhalu could increase the value of the firm and its stock price by leveraging the firm with perpetual debt. Accordingly, she perposed that Bhalu sell $ 2 million of 6% annual coupon bonds at par and use the proceeds to buy back some of its outstanding shares. the following information pertains to Bhalu before the repuerchase secision:

Value of the firm: $10 million

Outstanding shares: 500,000

Annual pre-tax earnings. $1.5 million

Tax rate: 40%

Determine the following:

a) Expected ROE(return on equity) prior to leveraging decision

b) Stock price prior to leveraging decision

c) the value of the firm immediately after the decision was announced

d) the number of shares that Bhalu purchased as a result of debt financing

e) the ROE after restructuring
Answer #1

a.

Annual pre tax earnings 1500000
Less: Tax @40% 600000
After tax profit 900000
Value of equity 10000000
ROE 9%

b. Stock price = 10,000,000 / 500,000 = $20

c. Weight of the debt = 2/ (10 -2) = 1/4

Value of the firm = ( 2million / 25%) - 2million = $6million

d. No of shares purchased = $2million / 20 = 100,000 shares

e)

Annual pre tax earnings 1500000
Les: Interest (2million *6%) 120000
EBT 1380000
Less: Tax @40% 552000
After tax profit 948000
Value of equity                                  6,000,000
ROE 16%